OpenStudy (anonymous):

Ok i have a graph with an amplitude of A which is 6. And the period is pi. It is a -sin graph. My guess for the equation of this graph was -6sin2x i tried this with parenthises and without no luck. Is there something else it could be or is this simply the wrong equation? Further description of graph: starts at orgin goes down and like a sin curve goes to -A then back up to A and back to zero. The period of this is pi.

jimthompson5910 (jim_thompson5910):

Can you post the actual graph?

OpenStudy (anonymous):

Nope i can't pull up the graph Swear what i say is right i've had two people look at it and screw up too.

jimthompson5910 (jim_thompson5910):

So your graph looks like this 

OpenStudy (anonymous):

Exactly i have no clue how you did that but yup thats it

OpenStudy (anonymous):

Thanks fot the help

jimthompson5910 (jim_thompson5910):

I used the draw feature

jimthompson5910 (jim_thompson5910):

The answer is y = -6*sin(2x), but you can easily do phase shifts to get equivalent answers. However, y = -6*sin(2x) is the most basic one.

jimthompson5910 (jim_thompson5910):

So there might be a typo somewhere

jimthompson5910 (jim_thompson5910):

could it be that the distance from peak to valley is 6 units? if so, then that would make the amplitude to be 3

OpenStudy (anonymous):

Thats what i put in but nothing. If you can think of anything else it would be appreciated but you helped the most for sure

jimthompson5910 (jim_thompson5910):

So you also tried y = -3*sin(2x) and that didn't work?

OpenStudy (anonymous):

I'll try that too thanks

jimthompson5910 (jim_thompson5910):

let me know how it goes

OpenStudy (anonymous):

No luck locked out but i had no shot anyway thanks

jimthompson5910 (jim_thompson5910):

np